# A Survey on Deep Learning Hardware Accelerators for Heterogeneous HPC Platforms

### v2

@@ 5.4. Accelerators based on Multi-Chip Modules
The alternate multichip-module (MCM) silicon interposer-based integration technology, described in Section A.9 of Appendix A, offers several advantages over single-chip designs. These advantages include increased functionality, reduced power consumption, higher performance, improved reliability, and cost savings. By utilizing MCM, designers can combine multiple chips and functionalities into a single package, resulting in a reduced overall footprint and cost. MCM-based designs can also be optimized for different power requirements, achieving higher power efficiency compared to single-chip designs. Additionally, MCM-based designs can be fine-tuned for different tasks, offering superior performance compared to single-chip counterparts.
The use of redundancy in MCM-based designs improves fault tolerance, leading to improved reliability compared to single-chip designs. Furthermore, MCM-based designs can utilize off-the-shelf components and existing manufacturing processes and technology, contributing to cost savings in overall manufacturing.
@@
rounded/.style=ellipse,draw, squared/.style=rectangle,draw, qtree, [Accelerators based on Multi-Chip Modules [General Purpose [ Mounce et al. (mounce_ac16, )
Vijayaraghavan et al. (vijayaraghavan_hpca17, )
Arunkumar et al. (arunkumar_isca17, ), Lin et al. (lin_vlsi19, )
Vivet et al. (vivet_jssc21, )
Martinez et al. (martinez_vlsi20, ) ] ], [Domain-Specific [ Kwon et al. (kwon_iceic23, ), Nurvitadhi et al. (nurvitadhi_fccm19, )
Verhelst et al. (verhelst2022ml, ), Arvon (tang_jssc24, )
Lan et al. (lan_eptc21, ), Lego (xuan_isocc22, )
Chimera (prabhu2022chimera, ), SWAP (sharma2022swap, )
SPRINT (li2021sprint, ), Simba (zimmer_jssc20, ; shao_micro19, )
Instinct (smith_isscc24, ), DOJO (talpes_micro23, ) ] ] ]
@@
In the realm of DL, chiplet-based design is utilized to create hardware accelerator platforms that are both efficient and scalable. Designing AI processors for data explosion computing due to the physical limitations of semiconductors and high costs is challenging. In (kwon_iceic23, ) authors propose chiplet-based design as a viable solution to this problem. They outline various aspects of designing a chiplet AI processor, including incorporating NPU chiplets, HBM chiplets, and 2.5D interposers, ensuring signal integrity for high-speed interconnections, power delivery network for chiplets, bonding reliability, thermal stability, and interchiplet data transfer on heterogeneous integration architecture. They conclude that chiplet-based design provides higher performance at a lower cost compared to IP-based design.
@@
Data-intensive DL algorithms with strict latency constraints require balancing both data movement and compute capabilities. Thus, persistent approaches that keep the entire DL model on-chip are becoming the new norm for real-time services to avoid expensive off-chip memory accesses. In (nurvitadhi_fccm19, ) it is shown how the integration of FPGA with ASIC chiplets outperforms GPU-based platforms (NVIDIA Volta) in terms of latency by enhancing on-chip memory capacity and bandwidth and provides compute throughput matching the required bandwidth. Specifically, it is reported that the GPU and chiplet-based FPGA computing capabilities are 6% and 57% of their peak, respectively. In terms of delay and energy efficiency, the FPGA outperforms the GPU with a delay that is 1/16 and energy efficiency that is 34x better than the GPU’s peak performance.
@@
In (verhelst2022ml, ) the authors investigate the recent multi-core trend in deep-learning accelerators evolution as a solution to further increase throughput and match the ever-growing computational demands. Chiplet integration is considered a promising implementation strategy for both homogeneous and heterogeneous multi-core accelerators. The trend towards developing high throughput and energy-efficient neural network hardware accelerators due to the growing complexity and dimension of neural network algorithms is analyzed in (lan_eptc21, ). The authors propose a chiplet-based architecture for a multi-core neuromorphic processor with a chip-package co-design flow. It is shown how the proposed design is reusable for different neuromorphic computing applications by scaling the number of chips in a package and by reusing existing IPs from different technology nodes with 2.5D integration technology. The challenges of using modern DNN accelerators in multi-tenant DNN data centers are investigated in (xuan_isocc22, ). The MCM architecture is proposed as a promising approach to address this issue but highlights the challenge of distributing DNN model layers with different parameters across chiplets. Thus the authors present Lego MCM architecture with a dynamic scheduler that adapts to the size of DNN model layers and increases chiplet utilization. The results show that Lego MCM achieves a 1.51x speedup over a monolithic DNN accelerator. Chimera (prabhu2022chimera, ) is a non-volatile chip for DNN training and inference that does not require off-chip memory. Multiple Chimera accelerator chiplets can be combined in a multi-chip system to enable inference on models larger than the single-chip memory with only $5\%$ energy overhead. The Arvon accelerator (tang_jssc24, ) is a heterogeneous System-in-Package that integrates FPGA and DSP chiplets designed for workload acceleration. Key characteristics include heterogeneous integration, combining a 14-nm FPGA chiplet with two 22-nm DSP chiplets using embedded multidie interconnect bridge technology. The Arvon platform is designed to support various workloads, including neural network processing through a specific compilation procedure that optimizes workload distribution across the FPGA and DSP chiplets.
@@
Notable examples of commercial chiplet-based hardware accelerators include AMD’s Instinct (smith_isscc24, ) and Tesla’s DOJO (talpes_micro23, ). The AMD Instinct MI300 series accelerators are designed for HPC and AI at exascale levels. They utilize a modular chiplet architecture that integrates data center-class CPUs, GPU accelerated compute, AMD Infinity Cache, and 8-stack HBM3 memory in a single package. Specifically, the MI300X model targets traditional dual-processor CPU servers with eight GPU accelerators, host DDR, and device HBM for large AI model training and inference. The MI300A model combines three Zen 4 CPU chiplets with six CDNA 3 GPU chiplets for high-density HPC systems, facilitating seamless CPU-GPU interaction without explicit data transfers. The DOJO system, developed by Tesla, is an exascale computer designed for machine learning training applications. It features a homogeneous modular architecture with a hierarchical chiplet-based organization. Each chiplet, named D1 die, contains 354 DOJO nodes, each functioning as a full-fledged computer with dedicated processing pipelines, local memory, and network interfaces. A training tile is formed by integrating 25 D1 dies within a single package.
@@
It is worth noting that the inter-chiplet communication system plays a crucial role as it significantly affects the overall energy efficiency and performance metrics of the system. In this direction, in SWAP (sharma2022swap, ), a DNN inference accelerator based on the 2.5D integration of multiple resistive RAM chiplets, a design space exploration flow is proposed to optimize the interconnection Network-on-Package, minimizing inter-chiplet communications and enabling link pruning. In SPRINT (li2021sprint, ), a photonic-based interconnects is used as an alternative to metallic-based inter-chiplet networks for the design of a DNN inference hardware accelerator.
@@ 5.5. Accelerators based on Quantum Computing and Photonic Computing
There is a general agreement that Quantum computers will not replace conventional computing systems, but they will be used in combination with supercomputers to accelerate some hard-to-compute problems. Quantum computers will play the role of unconventional accelerators to outperform conventional supercomputers, thanks to the improved parallelism that enables the so-called quantum speedup. Governments, supercomputing centers, and companies around the world have also started to investigate $How$/$When$/$Where$ quantum processing units (QPUs) could fit into HPC infrastructures to speed up some heavy tasks, such as Deep Learning workloads. Emerging trends and commercial solutions related to hybrid quantum-classical supercomputers are described in  (HPCWIRE2022, ). To address this challenging trend, in October 2022, the EuroHPC Joint Undertaking initiative selected six supercomputing centers across the European Union to host quantum computers and simulators. IBM Research was the first provider to offer a cloud-based QC service. IBM Qiskit  (QISKIT2023, ) is an open-source SDK based on a library of quantum gates/circuits: Remote users can develop quantum programs and execute them on quantum simulators and cloud-based quantum processors. Cloud providers have also jumped into the quantum race. As an example, Amazon Braket (BRAKET2023, ) is a QC service based on different types of quantum systems and simulators, including the quantum annealer from D-Wave. On this trend, there is a general agreement that GPUs will play a key role in hybrid quantum-classical computing systems. GPU company NVIDIA offers CuQuantum DGX hardware appliance integrating a software container on a full-stack quantum circuit simulator: The system uses NVIDIA’s A100 GPUs to accelerate quantum simulation workloads.
@@
Recently, a survey on QC technologies appeared in  (GYONGYOSI201951, ), while another survey on QC frameworks appeared in  (upama2022, ). More specifically, there is a promising research trend on Quantum Machine Learning  (Biamonte2017, ) which aims at developing quantum algorithms that outperform classical computing algorithms on machine learning tasks such as recommendation systems. More in detail, classical DNNs inspired the development of Deep Quantum Learning methods. The main advantage of these methods is that they do not require a large, general-purpose quantum computer. Quantum annealers, such as the D-Wave commercial solutions (D-WAVE2023, ), are well-suited for implementing deep quantum learners. Quantum annealers are special-purpose quantum processors that are significantly easier to construct and scale up than general-purpose quantum computers. Following this research trend, Google proposed TensorFlow Quantum (TFQ)  (broughton2021, ), an open-source quantum machine learning library for prototyping hybrid quantum-classical ML models.
@@
The second challenging and promising research direction is represented by the use of Photonic Computing to further accelerate DL tasks. Photonic Computing relies on the computation of electromagnetic waves typically via non-linear modulation and interference effects. It was originally introduced in the 1980s to address optical pattern recognition and optical Fourier transform processing (ambs_2010, ). Despite the potential advantages of processing parallelism and speed, optical computing has never translated into a widely adopted commercial technology. Only recently, due to the emergence of data-intensive computing tasks, such as AI, optical computing has seen a renewed interest. There are two main advantages of optical computing, namely (i) the inherent speed of signal transmission, where light pulses can be transferred without the typical RC delays and IR drop of electrical interconnects, and (ii) the inherent parallelism, where multiple wavelengths, polarizations, and modes can be processed by the same hardware (e.g., waveguides, interferometers, etc.), without interfering with each other. These properties can provide strong benefits to data-intensive computing tasks such as DL. Photonic computing represents a promising platform for accelerating AI. For instance, it has been estimated that photonic multiply-accumulate operations can show significant improvements over digital electronics in terms of energy efficiency ($>10^{2}$), speed ($>10^{3}$), and compute density ($>10^{2}$(nahmias_2020, ). However, there are still many challenges to developing an industrially feasible photonic system. The main challenge is the area/energy inefficiency of processing across the mixed optical/electronic domain. Optical-electrical conversion and vice versa result in considerable overhead in terms of area and power consumption. To bridge this gap, the trend is developing silicon photonic integrated circuits (PICs) with increasing robustness, manufacturability, and scalability. Photonic computing essentially operates in the analog domain, thus accuracy is deeply affected by accumulated noise and imprecision of optical devices, such as electro-optic and phase change modulators. These challenges, similar to those arising in analog IMC, might be mitigated by hardware-aware training and system-level calibration techniques.
@@ A.5. Accelerating Arithmetic Data-paths
Performance achievable with ASIC accelerators for the inference of DL models is mainly dependent on the structure of the arithmetic data path. At its core, DL systems perform several finite impulse response operations over a large set of data. Hence the accelerator can be optimized by exploiting techniques used for the efficient implementation of the underlying arithmetic operations. As shown in Figure  12, three main types of optimization can be performed on the arithmetic data path. The first optimization approach takes into account that convolution is one of the main operations performed in a DL system. As demonstrated in (Cheng_2004, ; Tsao_2012, ), mono-dimensional convolutions can be efficiently computed by performing a reduced number of multiplications, thus improving the trade-off between the throughput of the circuit and the number of hardware resources needed for its implementation. This technique has been further developed in (Wang_2018, ; Cheng_2020, ; Wang_2022, ), where it is applied to multi-dimensional convolutions.
@@
Figure 12. Taxonomy of the data-path architectures described in Section LABEL:sssec:datapath
@@ A.8. RRAM and PCM Technologies for IMC
Figure 14. (a) RRAM and PCM devices structure and (b) their arrangement in a crossbar structure for matrix-vector multiplication. (c) Example of a stand-alone DNN accelerator (i.e., PRIME (chi2016isca, )) using RRAM crossbars for in situ MAC operations. Reprinted from (lepri2023jeds, ) and (chen2020engineering, ) under Creative Commons License.
@@ A.9. Alternative Integration Technologies
Figure 15. Die-level integration through TSV-based 3D and interposer-based 2.5D technologies 15(a) (stow_iccad16, ) and overall number of chips and impact on yield of an example defect distribution for two different chip sizes (ajaykumar_micro15, )

### v3

@@ 5.4. Accelerators based on Multi-Chip Modules
The alternate multichip-module (MCM) silicon interposer-based integration technology, described in Section A.9 of Appendix A, offers several advantages over single-chip designs, including increased functionality, reduced power consumption, higher performance, improved reliability, and cost savings. By utilizing MCM, designers can combine multiple chips and functionalities into a single package, resulting in a reduced overall footprint and cost. Furthermore, MCM-based designs can utilize off-the-shelf components and existing manufacturing processes and technology, contributing to cost savings in overall manufacturing.
@@
In the realm of DL, chiplet-based design is utilized to create hardware accelerators that are both efficient and scalable. The chiplet-based design proposed in (kwon_iceic23, ) is a viable solution to provide higher performance at a lower cost compared to IP-based design. In (kwon_iceic23, ), various aspects of designing a chiplet AI processor are considered, including incorporating NPU chiplets, HBM chiplets, and 2.5D interposers, ensuring signal integrity for high-speed interconnections, power delivery network for chiplets, bonding reliability, thermal stability, and interchiplet data transfer on heterogeneous integration architecture.
@@
rounded/.style=ellipse,draw, squared/.style=rectangle,draw, qtree, [Accelerators based on Multi-Chip Modules [General Purpose [ Mounce (mounce_ac16, ), Vijayaraghavan (vijayaraghavan_hpca17, )
Arunkumar (arunkumar_isca17, ), Lin (lin_vlsi19, )
Vivet (vivet_jssc21, ), Martinez (martinez_vlsi20, ) ] ], [Domain-Specific [ Kwon (kwon_iceic23, ), Nurvitadhi (nurvitadhi_fccm19, ), Verhelst (verhelst2022ml, )
Arvon (tang_jssc24, ), Lan (lan_eptc21, ), Lego (xuan_isocc22, ) Chimera (prabhu2022chimera, )
SWAP (sharma2022swap, ), SPRINT (li2021sprint, ), Simba (zimmer_jssc20, ; shao_micro19, ), Instinct (smith_isscc24, ), DOJO (talpes_micro23, ) ] ] ]
@@
At the aim of balancing both data movement and compute capabilities of data-intensive DL algorithms, keeping the entire DL model on-chip is becoming the new norm for real-time services to avoid expensive off-chip memory accesses. In (nurvitadhi_fccm19, ) it is shown how the integration of FPGA with ASIC chiplets enhances on-chip memory capacity and bandwidth and provides compute throughput that outperforms GPU-based platforms (NVIDIA Volta). Specifically, the GPU and chiplet-based FPGA computing capabilities are 6% and 57% of their peak, respectively. Moreover, the FPGA achieves a delay that is 1/16 and energy efficiency that is 34x better than the GPU.
@@
In accordance with the recent trend in DL accelerators, chiplet integration is considered a promising implementation strategy for both homogeneous and heterogeneous multi-core accelerators to further increase throughput and match the ever-growing computational demands  (verhelst2022ml, ). . In (lan_eptc21, ), a chiplet-based architecture is proposed for a multi-core neuromorphic processor with a chip-package co-design flow. The proposed design is reusable for different neuromorphic computing applications by scaling the number of chips in a package and by reusing existing IPs from different technology nodes with 2.5D integration technology. The MCM architecture presented in (xuan_isocc22, ) is a promising approach to address the issue of using modern DNN accelerators in multi-tenant DNN data centers, but it leaves the challenge of distributing DNN model layers with different parameters across chiplets still open. When a dynamic scheduler is used to comply with the size of DNN model layers and increase chiplet utilization, the Lego MCM architecture achieves a 1.51x speedup over a monolithic DNN accelerator. Chimera (prabhu2022chimera, ) is a non-volatile chip for DNN training and inference that does not require off-chip memory. Multiple Chimera accelerator chiplets can be combined in a multi-chip system to enable inference on models larger than the single-chip memory with only $5\%$ energy overhead. The Arvon accelerator (tang_jssc24, ) is a heterogeneous System-in-Package that integrates a 14-nm FPGA chiplet with two 22-nm DSP chiplets using embedded multidie interconnect bridge technology. Arvon is designed to support various workloads, including neural network processing through a specific compilation procedure that optimizes workload distribution across the FPGA and DSP chiplets.
@@
Notable examples of commercial chiplet-based hardware accelerators include AMD’s Instinct (smith_isscc24, ) and Tesla’s DOJO (talpes_micro23, ). The AMD Instinct MI300 series are designed for HPC and AI at exascale levels. They utilize a modular chiplet architecture that integrates data center-class CPUs, GPU accelerated compute, AMD Infinity Cache, and 8-stack HBM3 memory in a single package. Specifically, the MI300X model targets traditional dual-processor CPU servers with eight GPU accelerators, host DDR, and device HBM for large AI model training and inference. The MI300A model combines three Zen 4 CPU chiplets with six CDNA 3 GPU chiplets for high-density HPC systems, facilitating seamless CPU-GPU interaction without explicit data transfers. The Tesla’s DOJO system is an exascale computer designed for ML training applications. It features a homogeneous modular architecture with a hierarchical chiplet-based organization. Each chiplet, named D1 die, contains 354 DOJO nodes, each functioning as a full-fledged computer with dedicated processing pipelines, local memory, and network interfaces. A training tile is formed by integrating 25 D1 dies within a single package.
@@
SWAP (sharma2022swap, ) is a DNN inference accelerator based on the 2.5D integration of multiple resistive RAM chiplets. In  (sharma2022swap, ), a design space exploration flow is proposed to optimize the interconnection Network-on-Package, minimizing inter-chiplet communications and enabling link pruning. Further improvements are achieved in SPRINT (li2021sprint, ), where a photonic-based interconnects is used as an alternative to metallic-based inter-chiplet networks.
@@ 5.5. Accelerators based on Quantum Computing and Photonic Computing
There is a general agreement that Quantum computers will not replace conventional computing systems, but they will be used in combination with supercomputers to accelerate some hard-to-compute problems. Quantum computers will play the role of unconventional accelerators to outperform conventional supercomputers, thanks to the improved parallelism that enables the so-called quantum speedup. Governments, supercomputing centers, and companies around the world have also started to investigate $How$/$When$/$Where$ quantum processing units (QPUs) could fit into HPC infrastructures to speed up some heavy tasks, such as DL workloads. Emerging trends and commercial solutions related to hybrid quantum-classical supercomputers are described in  (HPCWIRE2022, ). To address this challenging trend, in October 2022, the EuroHPC Joint Undertaking initiative selected six supercomputing centers across the European Union to host quantum computers and simulators. IBM Research was the first provider to offer a cloud-based QC service. IBM Qiskit  (QISKIT2023, ) is an open-source SDK based on a library of quantum gates/circuits: Remote users can develop quantum programs and execute them on quantum simulators and cloud-based quantum processors. Cloud providers have also jumped into the quantum race. As an example, Amazon Braket (BRAKET2023, ) is a QC service based on different types of quantum systems and simulators, including the quantum annealer from D-Wave. On this trend, there is a general agreement that GPUs will play a key role in hybrid quantum-classical computing systems. GPU company NVIDIA offers CuQuantum DGX hardware appliance integrating a software container on a full-stack quantum circuit simulator: The system uses NVIDIA’s A100 GPUs to accelerate quantum simulation workloads.
@@
Recently, a survey on QC technologies appeared in  (GYONGYOSI201951, ), while another survey on QC frameworks appeared in  (upama2022, ). More specifically, there is a promising research trend on Quantum Machine Learning  (Biamonte2017, ) which aims at developing quantum algorithms that outperform classical computing algorithms on ML tasks such as recommendation systems. More in detail, classical DNNs inspired the development of Deep Quantum Learning methods. The main advantage of these methods is that they do not require a large, general-purpose quantum computer. Quantum annealers, such as the D-Wave commercial solutions (D-WAVE2023, ), are well-suited for implementing deep quantum learners. Quantum annealers are special-purpose quantum processors that are significantly easier to construct and scale up than general-purpose quantum computers. Following this research trend, Google proposed TensorFlow Quantum (TFQ)  (broughton2021, ), an open-source quantum machine learning library for prototyping hybrid quantum-classical ML models.
@@
The second challenging and promising research direction is represented by the use of Photonic Computing to further accelerate DL tasks. Photonic Computing relies on the computation of electromagnetic waves typically via non-linear modulation and interference effects. It was originally introduced in the 1980s to address optical pattern recognition and optical Fourier transform processing (ambs_2010, ). Despite the potential advantages of processing parallelism and speed, optical computing has never translated into a widely adopted commercial technology. Only recently, due to the emergence of data-intensive computing tasks, such as AI, optical computing has seen a renewed interest. There are two main advantages of optical computing, namely (i) the inherent speed of signal transmission, where light pulses can be transferred without the typical RC delays and IR drop of electrical interconnects, and (ii) the inherent parallelism, where multiple wavelengths, polarizations, and modes can be processed by the same hardware (e.g., waveguides, interferometers, etc.), without interfering with each other. These properties can provide strong benefits to data-intensive computing tasks such as DL. Photonic computing represents a promising platform for accelerating AI. For instance, it has been estimated that photonic MAC operations can show significant improvements over digital electronics in terms of energy efficiency ($>10^{2}$), speed ($>10^{3}$), and compute density ($>10^{2}$(nahmias_2020, ). However, there are still many challenges to developing an industrially feasible photonic system. The main challenge is the area/energy inefficiency of processing across the mixed optical/electronic domain. Optical-electrical conversion and vice versa result in considerable overhead in terms of area and power consumption. To bridge this gap, the trend is developing silicon photonic integrated circuits (PICs) with increasing robustness, manufacturability, and scalability. Photonic computing essentially operates in the analog domain, thus accuracy is deeply affected by accumulated noise and imprecision of optical devices, such as electro-optic and phase change modulators. These challenges, similar to those arising in analog IMC, might be mitigated by hardware-aware training and system-level calibration techniques.
@@ A.5. Accelerating Arithmetic Data-paths
Performance achievable with ASIC accelerators for the inference of DL models is mainly dependent on the structure of the arithmetic data path. At its core, DL systems perform several finite impulse response operations over a large set of data. Hence the accelerator can be optimized by exploiting techniques used for the efficient implementation of the underlying arithmetic operations. As shown in Figure  11, three main types of optimization can be performed on the arithmetic data path. The first optimization approach takes into account that convolution is one of the main operations performed in a DL system. As demonstrated in (Cheng_2004, ; Tsao_2012, ), mono-dimensional convolutions can be efficiently computed by performing a reduced number of multiplications, thus improving the trade-off between the throughput of the circuit and the number of hardware resources needed for its implementation. This technique has been further developed in (Wang_2018, ; Cheng_2020, ; Wang_2022, ), where it is applied to multi-dimensional convolutions.
@@
Figure 11. Taxonomy of the data-path architectures.
@@ A.8. RRAM and PCM Technologies for IMC
Figure 13. (a) RRAM and PCM devices structure and (b) their arrangement in a crossbar structure for matrix-vector multiplication. Adapted with permission from (lepri2023jeds, ) under Creative Commons License (CC-BY-4.0).
@@ A.9. Alternative Integration Technologies
Figure 14. Die-level integration through TSV-based 3D and interposer-based 2.5D technologies 14(a) (stow_iccad16, ) and overall number of chips and impact on yield of an example defect distribution for two different chip sizes (ajaykumar_micro15, )